Please help me to solve this question =/
The sum of Ken's LDL and HDL cholesterol is 210, and his LDL to HDL ratio is 2.75. What are his LDL and HDL?

Answer by Alan3354(69443)

The sum of Ken's LDL and HDL cholesterol is 210, and his LDL to HDL ratio is 2.75. What are his LDL and HDL?
-------------
The ratio is 2.75:1 = 11:4
11+4 = 15
210/15 = 14
---------
LDL = 11*14 = 154
HDL = 4*4 = 56